OpenStudy (anonymous):

What is the 23rd term of the arithmetic sequence where a1 = 8 and a9 = 48 ?

OpenStudy (callisto):

Difference between each consecutive term = (48-8)/8 = 5 T(23) = 8 + (23-1)x5 = 8+110 = 118